# Symplectic Approach to Global Stability

### v1

@@ Abstract
Global Stability Symplectic Geometry Liouville-Mineur-Arnold Theorem Bi-Hamiltonian Systems.
@@ 2 Global Stability for a Vector Field
Let $M$ be a finite-dimensional smooth manifold in the sense of [14]. Assume that the dynamics is given by a vector field $X$ on $M$. We restrict attention to two physically relevant global stability notions [7], which can be defined in terms of the integrable curves $\gamma$ of $X$. We denote by $I_{\gamma}\subset\mathbb{R}$ the interval of definition of $\gamma$. We abide by the convention that a set $A$ is relatively compact with respect to a topological space $N$ if $A$ is contained in a compact set of $N$.
@@
if the image $Im(\gamma)$ is relatively compact with respect to $M$ for all integrable curves $\gamma$. This is also known as Lagrange stability.
@@
if, for all integrable curves $\gamma$, and for all subsets $S\subset I_{\gamma}$ relatively compact with to $\mathbb{R}$, its image $\gamma(S)$ is relatively compact with respect to $M$. This corresponds to $X$ being a complete vector field.
A non-complete vector field has an integrable curve that cannot be defined for all times. This implies that the image of the curve cannot be contained in any compact set of $M$ [14]. Then, we say that $X$ has a blow up. In particular, $G1$ implies $G2$.
@@
Let $f^{\neg}(\mu):=\{x\in M|f(x)=\mu\}\,$. If $M$ is Hausdorff, a proper map $f:M\rightarrow N$ is confining [7]. The converse is not true. For instance, $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ given by $f(x,y)=\sin(x^{2}+y^{2})$ is confining but not a proper map. If $M$ and $N$ are manifolds, $f$ is a smooth function and $\mu$ is a regular point of $N$, then the level set $f^{\neg}(\mu)$ is a closed manifold. As such, its connected components are path-connected. We have the following theorem (adapted from [7]):
@@
The condition $X(Q)=0$ means that $Q$ is a conserved quantity of $X$. Therefore, each integral curve of $X$ lies inside one path-connected component of a level set of $Q$, all of curveswhich are relatively compact with respect to $M$.
@@ 3 Extension to Several Symmetries
Let $M$ be a manifold with dimension $2n$. Consider the Hamiltonian system $(\omega,H)$, where $\omega\in\Omega^{2}(M)$ is a symplectic form and $H$ is a smooth function. Let $X_{H}$ be the Hamiltonian vector field associated to $(\omega,H)$. In particular, $H$ is a conserved quantity.
@@ 3.1 $G1$ Stability for Systems with Multiple Conserved Quantities
Therefore, each integral curves of $X_{H}$ is inside one path-connected component of a level set of $J$. Hence, they are relatively compact with respect to $M$.
Another possibility is that a functional combination of the conserved quantities is confining. As it turns out, if the momentum map $J$ is not confining (proper), then no functional combination of the conserved quantities is confining (proper).
@@
If $f$ is not confining, then there exists a point $c\in N$ such that a path-connected component $F$ of $f^{\neg}(c)$ cannot be contained in any compact set of $M$. Since $c\in g^{\neg}(g(c))$, we have that $(g\circ f)^{\neg}(g(c))\supset f^{\neg}(c)$. In particular, it has a path connected component $G\subset(g\circ f)^{\neg}(g(c))$ that contains $F$. Thus, $G$ cannot be contained in any compact set of $M$.
@@
If $f$ is not a proper map, then there exists a compact set $K\subset N$ such that $f^{\neg}(K)$ is not compact, even though it is closed because $f$ is continuous and $N$ is Hausdorff. Since $g$ is continuous, $g(K)$ is compact. Moreover, $(g\circ f)^{\neg}(g(K))\supset f^{\neg}(K)$. If $(g\circ f)^{\neg}(g(K))$ were compact, then $f^{\neg}(K)$ would be compact, which is a contradiction.
The converse of lemma 2 holds true in the following case. Assume that the Hamiltonian system $(\omega,H)$ has conserved quantities $\{Q_{i}\}_{i=1,\dots,r\leq n}$. Set $J:M\rightarrow\mathbb{R}^{r}$ as $J=(Q_{1},\dots,Q_{r})$ and define the function $P:M\rightarrow\mathbb{R}$ as $P:=\sum_{i=1}^{r}Q_{i}^{2}$.
@@
$J=(Q_{1},\dots,Q_{r})$ is a proper map iff $P$ is a proper map.
@@
Let $I\subset\mathbb{R}$ be a compact set with maximum $b\in\mathbb{R}$. Define $K:=[-\sqrt{|b|},\sqrt{|b|}]$ and let $K^{r}$ denote the Cartesian product of $r$ copies of $K$. Then, $Q_{i}(P^{\neg}(I))\subset K$ for all $i$. Therefore, $P^{\neg}(I)\subset\bigcap_{i=1}^{r}Q_{i}^{\neg}(K)$. Moreover,
$J^{\neg}(K^{r})=(Q_{1},\dots,Q_{r})^{\neg}(K^{r})=\bigcap_{i=1}^{r}Q_{i}^{\neg}(K)\,$ (2)
is a compact set because $J$ is a proper map and $K^{r}$ is a compact set. Therefore, $P^{\neg}(I)$ is inside a compact set. Besides, it is closed because $P$ is a continuous function. Thus, $P^{\neg}(I)$ is a compact set.
@@ 3.2 Global Stability for Integrable Systems
Consider a Hamiltonian system $(\omega,H)$ with $n$ conserved quantities $\{Q_{i}\}_{i=1,\dots n}$ in involution $\{Q_{i},Q_{j}\}=0$ such that $\mathrm{d}Q_{i}$ are generically independent. This implies that $H$ is a functional combination of the $Q_{i}$’s. The corresponding Hamiltonian vector fields, or symmetries, are defined as
@@
First, $F$ is confining. This is the standard Liouville-Mineur-Arnold theorem case. The connected components of generic level sets of $F$ are tori and action-angle variables are defined on a neighbourhood of these tori. Thus, the integrable curves are trapped inside tori and the system is $G1$ stable.
@@
If $\bar{\omega}_{A}(I_{a},\bar{I}_{A})$ is identically zero for all $A$ and cylinder, then the system does not move along the $\mathbb{R}^{k}$ directions. That is, the dynamics is confined to the tori $\mathbb{T}^{n-k}$ inside each cylinder. Therefore, the system is $G1$ stable on the cylinders. Moreover, in this situation, $s^{A}$ are also conserved quantities. Furthermore, the map
@@ 4 An application to physics: ghost stabilization
A vector field $X$ is ghost-ridden if, for all symplectic structures $(\omega,H)$ such that $i_{X}\omega=\mathrm{d}H$, $H$ is neither bounded from above nor from below.
@@ 5 Summary and outlook
We have provided a more exhaustive analysis in the fortunate case that, on a $2n$-dimensional symplectic manifold, at least $n$ independent conserved quantities in involution are known. We have shown that, for complete symmetries on the level sets of regular points, the integrable curves are inside a compact set if, and only if, there exists an invariant confining map. If all points are regular and symmetries are complete, then the result holds globally, see theorem 3.1.
@@
Thus far, global stability proofs hinge on the existence of conserved quantities with well-disposed properties. However, this is not required by definition. For three examples conjectured to be globally stable without any confining conserved quantity, see [6]. The development of a technique to prove global stability without appealing for conserved quantities would entail a paradigm shift in ghost-ridden systems and comprises a compelling line of future research.

### v2

@@ Abstract
Global Stability Symplectic Geometry Liouville-Mineur-Arnold Theorem.
@@ 2 Global Stability for a Vector Field
Let $M$ be a finite-dimensional smooth manifold in the sense of [14]. Assume that the dynamics is given by a vector field $X$ on $M$. We restrict attention to two physically relevant global stability notions [7], which can be defined in terms of the integral curves $\gamma$ of $X$. We denote by $I_{\gamma}\subset\mathbb{R}$ the interval of definition of $\gamma$. We abide by the convention that a set $A$ is relatively compact with respect to a topological space $N$ if $A$ is contained in a compact set of $N$.
@@
if the image $Im(\gamma)$ is relatively compact with respect to $M$ for all integral curves $\gamma$. This is also known as Lagrange stability.
@@
if, for all integral curves $\gamma$, and for all subsets $S\subset I_{\gamma}$ relatively compact with respect to $\mathbb{R}$, its image $\gamma(S)$ is relatively compact with respect to $M$. This corresponds to $X$ being a complete vector field.
A non-complete vector field has an integral curve that cannot be defined for all times. This implies that the image of the curve cannot be contained in any compact set of $M$ [14]. Then, we say that $X$ has a blow up. In particular, $G1$ implies $G2$.
To illustrate these definitions, consider the manifold $\mathbb{R}^{2}$ with coordinates $(q,v)$, and the vector fields:
$X_{1}=v\frac{\partial}{\partial q}-q\frac{\partial}{\partial v},\qquad X_{2}=v\frac{\partial}{\partial q},\qquad X_{3}=v\frac{\partial}{\partial q}+q^{2}\frac{\partial}{\partial v}\,.$ (1)
The first vector field $X_{1}$ corresponds to the harmonic oscillator and it is $G1$ stable. The second vector field $X_{2}$ corresponds to the free particle, and it is $G2$ stable but not $G1$. The last vector field $X_{3}$ is not $G2$ stable because it has at least an integral curve $\gamma(t)=(6t^{-2},-12t^{-3})$ that cannot be defined for all times. In references [5, 6, 7], the reader can find additional families of dynamical systems whose global stability has been determined.
@@
Let $f^{{-1}}(\mu):=\{x\in M|f(x)=\mu\}\,$. If $M$ is Hausdorff, a proper map $f:M\rightarrow N$ is confining [7]. The converse is not true. For instance, $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ given by $f(x,y)=\sin(x^{2}+y^{2})$ is confining but not a proper map. If $M$ and $N$ are manifolds, $f$ is a smooth function and $\mu$ is a regular point of $N$, then the level set $f^{{-1}}(\mu)$ is a closed manifold. As such, its connected components are path-connected. We have the following theorem (adapted from [7]):
@@
The condition $X(Q)=0$ means that $Q$ is a conserved quantity of $X$. Therefore, each integral curve of $X$ lies inside one path-connected component of a level set of $Q$, all of which are relatively compact with respect to $M$. Thus, all integral curves of $X$ are relatively compact with respect to $M$.
@@ 3 Extension to Several Symmetries
Let $M$ be a manifold with dimension $2n$. Consider the Hamiltonian system $(\omega,H)$, where $\omega\in\Omega^{2}(M)$ is a symplectic form and $H$ is a smooth function. Let $X_{H}$ be the Hamiltonian vector field associated to $(\omega,H)$, that is [1]
$i_{X_{H}}\omega=\mathrm{d}H\,.$ (2)
In particular, $H$ is a conserved quantity.
@@ 3.1 $G1$ Stability for Systems with Multiple Conserved Quantities
Therefore, each integral curve of $X_{H}$ is inside one path-connected component of a level set of $J$. Hence, they are all relatively compact with respect to $M$.
Given $r$ functions $Q_{i}:M\rightarrow\mathbb{R}$, for $i=1,\dots,r$ on $M$, and a map $g:\mathbb{R}^{r}\rightarrow V$, the functional combination of $\{Q_{i}\}_{i=1,\dots,r}$ by $g$ is the map $g(Q_{1},\dots,Q_{r}):M\rightarrow V$. If $Q_{i}$ are conserved quantities, $V$ is a manifold and $g$ is also at least once differentiable, then $g(Q_{1},\dots,Q_{r})$ is another conserved quantity.
Consequently, it is reasonable to ask if a functional combination of the conserved quantities is confining. As it turns out, if the momentum map $J$ is not confining (proper), then no functional combination of the conserved quantities is confining (proper).
@@
If $f$ is not confining, then there exists a point $c\in N$ such that a path-connected component $F$ of $f^{{-1}}(c)$ cannot be contained in any compact set of $M$. Since $c\in g^{{-1}}(g(c))$, we have that $(g\circ f)^{{-1}}(g(c))\supset f^{{-1}}(c)$. In particular, it has a path connected component $G\subset(g\circ f)^{{-1}}(g(c))$ that contains $F$. Thus, $G$ cannot be contained in any compact set of $M$.
@@
If $f$ is not a proper map, then there exists a compact set $K\subset N$ such that $f^{{-1}}(K)$ is not compact, even though it is closed because $f$ is continuous and $N$ is Hausdorff. Since $g$ is continuous, $g(K)$ is compact. Moreover, $(g\circ f)^{-1}(g(K))\supset f^{{-1}}(K)$. If $(g\circ f)^{-1}(g(K))$ were compact, then $f^{{-1}}(K)$ would be compact, which is a contradiction.
The converse of lemma 2 holds true in the following case. Assume that the Hamiltonian system $(\omega,H)$ has conserved quantities $\{Q_{i}\}_{i=1,\dots,r\leq n}$. Set $J:M\rightarrow\mathbb{R}^{r}$ as $J=(Q_{1},\dots,Q_{r})$ and define the function $P:M\rightarrow\mathbb{R}$ as $P:=\displaystyle\sum_{i=1}^{r}Q_{i}^{2}$.
@@
The map $J=(Q_{1},\dots,Q_{r})$ is a proper map if, and only if, $P$ is a proper map.
@@
Let $I\subset\mathbb{R}$ be a compact set with maximum $b\in\mathbb{R}$. Define $K:=[-\sqrt{|b|},\sqrt{|b|}]$ and let $K^{r}$ denote the Cartesian product of $r$ copies of $K$. Then, $Q_{i}(P^{{-1}}(I))\subset K$ for all $i$. Therefore, $P^{{-1}}(I)\subset\bigcap_{i=1}^{r}Q_{i}^{{-1}}(K)$. Moreover,
$J^{{-1}}(K^{r})=(Q_{1},\dots,Q_{r})^{{-1}}(K^{r})=\bigcap_{i=1}^{r}Q_{i}^{{-1}}(K)\,$ (4)
is a compact set because $J$ is a proper map and $K^{r}$ is a compact set. Therefore, $P^{{-1}}(I)$ is inside a compact set. Besides, it is closed because $P$ is a continuous function. Thus, $P^{{-1}}(I)$ is a compact set.
@@ 3.2 Global Stability for Integrable Systems
[…]
@@
First, $F$ is confining. This is the standard Liouville-Mineur-Arnold theorem case. The connected components of generic level sets of $F$ are tori and action-angle variables are defined on a neighbourhood of these tori. Thus, the integral curves are trapped inside tori and the system is $G1$ stable.
@@
If $\bar{\omega}_{A}(I_{a},\bar{I}_{A})$ is identically zero for all $A$ and every cylinder, then the system does not move along the $\mathbb{R}^{k}$ directions. That is, the dynamics is confined to the tori $\mathbb{T}^{n-k}$ inside each cylinder. Therefore, the system is $G1$ stable on the cylinders. Moreover, in this situation, $s^{A}$ are also conserved quantities. Furthermore, the map
@@ 4 An application to physics: ghost stabilization
A vector field $X$ on a manifold $M$ is ghost-ridden if, for all Hamiltonian systems $(\omega,H)$ on a manifold $M$ such that $i_{X}\omega=\mathrm{d}H$, $H$ is neither bounded from above nor from below.
@@ 5 Summary and outlook
We have provided a more exhaustive analysis in the fortunate case that, on a $2n$-dimensional symplectic manifold, at least $n$ independent conserved quantities in involution are known. We have shown that, for complete symmetries on the level sets of regular points, the integral curves are inside a compact set if, and only if, there exists an invariant confining map. If all points are regular and symmetries are complete, then the result holds globally, see theorem 3.1.
@@
Thus far, global stability proofs hinge on the existence of conserved quantities with well-disposed properties. However, this is not required by definition. For three examples for which there exists numerical evidence in favor of global stability without any confining conserved quantity, see [6]. The development of a technique to prove global stability without appealing for conserved quantities would entail a paradigm shift in ghost-ridden systems and comprises a compelling line of future research.